ニュービジネス特論 2 単位 (選択)

Management Theory of New Business  教務委員会副委員長 / 工学部，第一線の実務経験者 · 非常勤講師

【授業目的】ニュービジネスとは，新しいアイデアや専門的な知識 · 技術を駆使 ; して創造的に展開される事業を意味する言葉であり，その主たる担い手はベ ; ンチャーと呼ばれる企業である．この授業の目的は，受講生がベンチャー企 ; 業を起業するために必要とされる知識，ノウハウ，そしてスピリットを提供 ; することにある． ; 【授業概要】活力ある日本社会の再生が求められている．こうした要請を受け， ; 政府は平成 14 から 16 年度にかけて「大学発ベンチャー 3 年 1000 社計画」を ; 実施し，その目標はほぼ達成されたが，今後も継続的に起業家教育を推進し ; ていくことの重要性には変わりはない．この授業は，こうした認識にもとづ ; いて，徳島県及び財団法人とくしま産業振興機構が支援して開設された『と ; くしま経営塾「平成長久館」起業力養成講座』である． ; 【授業形式】 [形態] ; 【キーワード】 [キーワード] ; 【先行科目】 [先行科目] ; 【関連科目】 [関連科目] ; 【履修要件】学部で「ニュービジネス概論」の単位を修得していない者のうち，授 ; 業開始時間までに教室入りできる者，グループを編成しビジネスプランを作 ; 成する意思がある者に限り受講を認める． ; 【履修上の注意】遅刻や授業中の私語は厳禁とする．悪質な場合は，退場処分とし， ; 以後の受講を認めない場合もあるので，くれぐれも注意すること．講師の都 ; 合により，多少変更の可能性がある． ; 【到達目標】ベンチャービジネスを起業するために必要な知識を習得するととも ; に，ビジネスプランを作成できるようになることを目標とする． ; 【授業計画】 ; 1. ガイダンス ; 2. ニュービジネスとは? ; 3. 基調講演「二坪ビジネス成功の秘訣」 (仮題) ; 4. 独立型ベンチャー成功のための理論 ; 5. 起業者に必要な法知識 ; 6. 資金調達と資本政策 ; 7. 間接金融 ; 8. 直接金融 ; 9. 会社経営の基礎 10. 企業会計の基礎知識

11. ビジネスプラン作成のポイント

12. 経営戦略とマーケティング

13. 製品開発と知的財産権

14. ビジネスプラン作成実習

15. 筆記試験

16. ビジネスプラン発表会

【成績評価基準】到達目標の達成度を筆記試験 (60%) とビジネスプランの提出 (40%) で評価し， 60%以上の評点を合格とする．成績は評点を 100 点満点に換算して算定する．また，筆記試験およびビジネスプランの提出は，開講された授業の 3 分の 2 以上に出席した受講生にのみ認めるものとする．

【教科書】毎回レジュメを配付する．【参考書】授業時間に数冊紹介する．【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216805 【対象学生】大学院に在籍するすべての学生のうち，本学で開講されたニュービジネス概論の単位をまだ修得していない者を対象とする．受講希望者多数の場合は履修者数を抽選等の方法で制限する場合もある．詳細については，受講上の注意を示した掲示を行うので，受講希望者は必ずその指示にしたがうこと．【連絡先】

⇒ 教務委員会副委員長【備考】この授業では，毎回第一線で活躍されている実務家やアドバイザーを講師として招き，専門家の立場から理論プラス実践に役立つ知識の提供を行う．また，この授業は主に「起業」のための知識を提供するものであるが，その内容は一般企業や研究所，自治体などへの就職をめざす者にとっても十分役立つものとなっている．

Management Theory of New Business 2 units (selection)

Vice chairperson of School Affairs Committee , Faculty of Engineering / Faculty of Engineering First-line men with experience of practical business · Part-time Lecturer

Target⟩ New business is the word which means the business to make the most of technical knowledge, technology, the new idea and which is developed creatively, and that main bearer is the enterprise called a venture. There is a purpose of this class in providing the knowledge which is necessary, and know-how and a spirit so that a taking lecture student may establish a venture business.

Outline⟩ 活力ある日本社会の再生が求められている．こうした要請を受け，政府は平成 14 から 16 年度にかけて「大学発ベンチャー 3 年 1000 社計画」を実施し，その目標はほぼ達成されたが，今後も継続的に起業家教育を推進していくことの重要性には変わりはない．この授業は，こうした認識にもとづいて，徳島県及び財団法人とくしま産業振興機構が支援して開設された『とくしま経営塾「平成長久館」起業力養成講座』である．

Style⟩ [形態] �������⟩ [キーワード] Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ [関連科目] Requirement⟩ The person whom the one with the classroom is constructed by

the class start time in the person who doesn’t learn the unit of ”Introduction to new business” in the college, and a group are organized, and it is limited to the person with the intention to make a business plan, and taking lecture is recognized.

Notice⟩ 遅刻や授業中の私語は厳禁とする．悪質な場合は，退場処分とし，以後の受講を認めない場合もあるので，くれぐれも注意すること．講師の都合により，多少変更の可能性がある．

Goal⟩ Learning of the necessary knowledge to establish venture business,and the

ability which can make a business plan.

Schedule⟩

1. Guidance

2. What is the new business?

3. Keynote lecture

4. Theory of a stand-alone version venture success

5. The law knowledge which is necessary for entrepreneur

6. Fund-raising and capital policy 7. Indirect finance

8. Direct finance

9. The basis of the company management

10. The elementary knowledge of the business accounting

11. The point of the business plan preparation

12. Management strategy and marketing

13. Product development and intellectual property

14. Business plan preparation practice

15. Examination

16. Business plan announcement meeting Evaluation Criteria⟩ It is evaluated by the presentation (40%) of the written

examination (60%) and the business plan. And, the presentation of the written examination and the business plan is decided to be recognized only for the taking lecture student who attended it for 2/3 and more of the classes that it is opened.

Textbook⟩ The summary is distributed every time. Reference⟩ Several are introduced in the school hours.

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216805

Student⟩ he person who hasn’t learned the unit of the new business outline that it is opened with this learning in all the students enrolled in the graduate school yet is made the target. The number of completion people may be restricted in the method such as a lottery in case of a taking lecture applicant many. As for the details, an applicant for taking lecture is to follow those directions because the notice which showed attention in taking lecture is done.

Contact⟩

⇒ Vice chairperson of School Affairs Committee , Faculty of Engineering

Note⟩ The man of business that it is active in the forefront every time, and an adviser are invited as an instructor, and working knowledge is provided for theory plus practice in this class from the specialist’s position. This class is the thing which is useful enough even if it is taken to the person who aims at a general enterprise and a laboratory, the employment to the self-governing body and so on.